If f(x)= -4/5x+10 and g(x)= 2x+1, what is g(f(15)
Answer by Alan3354(69443)

If f(x)= -4/5x+10 and g(x)= 2x+1, what is g(f(15)
----------------
Is it f(x)= (-4/5)x+10 ?
Or f(x)= -4/(5x)+10 ?
Or f(x)= -4/(5x+10) ?
---------
what is g(f(15) --- parentheses don't match
For g(f(15)),
Sub 15 for x in f(x).
Then sub that in g(x).